OpenStudy (anonymous):

15. Which equation has no solution? (1 point) 8 – (5v + 3) = 5v – 5 3m – 6 = 5m + 7 – m 3w + 4 – w = 5w – 2(w – 2) 7y + 9 = 7y – 6

OpenStudy (anonymous):

7y+9=7y-6 7y=7y-15 0!=15 The answer is 7y+9=7y-6.